OpenStudy (anonymous):

A neighborhood swim club charges a one-time fee of $25 to join and $85 each month. If n is the number of months of membership and t is the total amount spent, then the variables are related by the equation t = 85n + 25.

OpenStudy (anonymous):

Lets see

OpenStudy (anonymous):

Which variable is the independent variable?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

A. t B. n C. There is no relationship between the variables.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

t is dependant on n

OpenStudy (anonymous):

n however is not

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

?

OpenStudy (anonymous):

B. n

OpenStudy (anonymous):

Do u get it?

OpenStudy (anonymous):

u said it wasn't N

OpenStudy (anonymous):

I said t depends on n so t cannot be the answer

OpenStudy (anonymous):

t is dependant

OpenStudy (anonymous):

n is independant. it does not rely on t

OpenStudy (anonymous):

lol yea

OpenStudy (anonymous):

I think N

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

yay!!!1

OpenStudy (anonymous):

Which is the equation of the line?

OpenStudy (anonymous):

plz new question each time. it just bothers me.

OpenStudy (anonymous):

A. 3x + y = 9 B. 3x – y = 9 C. x – y = 9 D. x + 3y = 9

OpenStudy (anonymous):

Oh ok